OpenStudy (anonymous):

FAN+MEDAL+TESTIMONIAL:):)!! How many arrangements are there of the word MATHEMATICS? How many of these start with the letter M? How many of the arrangements in part a have the T’s together?

OpenStudy (anonymous):

I think that this is a permutation problem. So given that there are 11 letters in the word math, I will do the following calculation

OpenStudy (anonymous):

11P11

OpenStudy (anonymous):

Without restrictions

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

hai

OpenStudy (anonymous):

Hi Sir:)

OpenStudy (anonymous):

Am I on the right track?

OpenStudy (anonymous):

ya know I'm stuck on this question right now too :P

OpenStudy (anonymous):

This is a university level math don't worryXD

OpenStudy (anonymous):

:O really? its on my 8th grade quiz

OpenStudy (anonymous):

Forgive me I am in a Canadian school.

OpenStudy (anonymous):

lol ok i will try to help tho

OpenStudy (anonymous):

gimee a sec

OpenStudy (anonymous):

Arrangement of letters of a word is a permutation question right?

OpenStudy (anonymous):

As opposed to combination

OpenStudy (anonymous):

Hi everyone:)

OpenStudy (anonymous):

Am I on the right track?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

How many arrangements are there of the word MATHEMATICS? Rule: Start with the factorial of the number of letters in the word. Then, for each indistinguishable letter in the word, divide by the factorial of the number of times that letter occurs in the word. "MATHEMATICS" is an 11-letter word. If all the letters were distinguishable like in "MATHEmatICS", the answer would be 11! = 39916800 However, there are 2 indistinguishable M's 2 indistinguishable A's 2 indistinguishable T's Thus, using the rule, we divide 11! by (2!)(2!)(2!) \[\frac{ 11! }{ 2!*2!*2! }=\frac{ 39916800 }{ 8 }=4989600\] How many of these start with the letter M? That amounts to finding all the distinguishable arrangements of the 10-letter "word" "ATHEMATICS" and putting an M in the beginning of each. "ATHEMATICS" is a 10-letter "word" and it contains 2 indistinguishable A's 2 indistinguishable T's Thus, using the rule, we divide 10! by (2!)(2!) \[\frac{ 10! }{ 2!*2! }=\frac{ 3628800 }{ 4 }=907200\] How many of the arrangements in part a have the T’s together? That amounts to finding all the distinguishable arrangements of the 10-letter "word" "MATHEMAICS" and inserting another T to the right of the "T" in each. "MATHEMAICS" is a 10-letter "word" and it contains 2 indistinguishable M's 2 indistinguishable A's Thus, using the rule, it's exactly the same answer as the second part. We divide 10! by (2!)(2!) \[\frac{ 10! }{ 2!*2! }=\frac{ 3628800 }{ 4 }=907200\] There ya go!

OpenStudy (anonymous):

YOU ROCK:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok testimonial?

OpenStudy (anonymous):

let me know if it could be better